find the area of the region in the first quadrant by bounded by the curve y=sqrt x^2+5, the x-axis and the line x=2


1
Expert's answer
2022-05-27T00:26:20-0400

ANSWER: Area of region "\cong 5.0118"

EXPLANATION:

Let "D=\left \{ (x,y):0\leq x\leq2,\, 0\leq y\leq \sqrt{x^{2}+5} \right \} , A=Area" of "D" , then

"A=\int_{0}^{2}\sqrt{x^{2}+5}dx"

We calculate the integral by parts : denote "u= \sqrt{x^{2}+5} ,\, dv=dx\," . Therefore , "v=x, \, du=\frac{2x}{2\sqrt{x^{2}+5}}=\frac{ x}{ \sqrt{x^{2}+5}}" and

"A=\left [ x\cdot \sqrt{x^{2}+5} \right ]_{0}^{2}-\int_{0}^{2}\frac{x^{2}}{\sqrt{x^{2}+5}}dx=\left [ 2\cdot \sqrt{2^{2}+5}-0 \right ] -\int_{0}^{2}\frac{x^{2}+5-5}{\sqrt{x^{2}+5}}dx=6-A+5\cdot \int_{0}^{2}\frac{1}{\sqrt{x^{2}+5}}dx=6-A+5\cdot \left [ \ln \left | x+\sqrt{x^{2}+5} \right | \right ]_{0}^{2}=6-A+5\cdot \left ( \ln (2+3) -\ln \sqrt{5}\right )=6-A+5\ln \frac{5}{\sqrt{5}}=6-A+5\cdot\frac{1}{2}\ln 5 ."

From the equality "A=6-A+5\cdot\frac{1}{2}\ln 5" it follows "A=3+\frac{5}{4}\ln 5 \cong 5.0118."